# L3Cube-MahaEmotions: A Marathi Emotion Recognition Dataset with Synthetic Annotations using CoTR prompting and Large Language Models

### v1

@@ 4.2 GPT-4 vs Llama3-405B
After annotation, we trained a classifier on the GPT-labeled dataset. The overall performance of this classifier on the test set is shown in Table 3. It achieved an accuracy of 63%, with a precision of 0.45, a recall of 0.50, and an F1 score of 0.47. The detailed confusion matrix is presented in Figure 3, which shows the classification behavior across all emotion categories.
The confusion matrix in Figure 3 shows that the classifier misclassifies various emotion classes as Neutral. For example, out of 68 total Surprise samples, 20 were predicted as Neutral. Similarly, for Excitement, 9 out of 45 instances were misclassified as Neutral, and in Anger/Disgust, where 48 samples were misclassified as Neutral.
These results point to the classifier’s difficulty in distinguishing subtle emotional expressions from truly neutral content. Many Marathi sentences carry emotional nuances that may not be overtly visible. This can lead to confusion with neutral statements. This issue is particularly critical in low-resource languages like Marathi, where emotional subtleties often rely on cultural or contextual cues rather than explicit linguistic markers. Therefore, improving classification of these subtle emotions is a key area for future work.
@@ 5 Future Work and Conclusion
In the future, we plan to improve this work by training LLMs on more Marathi-specific emotion data to better understand the language and its expressions. We also aim to include sentences which show complex emotions. Another important direction is testing our method on more LLMs, such as Gemma, Grok, DeepSeek, and Mistral, to check how well it works across different models. Lastly, our approach can be extended to other Indian languages that also lack emotion datasets, helping build better NLP tools for a wider range of users.
This approach can further be used for other low-resource Indic languages.

### v2

@@ 4.2 GPT-4 vs Llama3-405B
After annotation, we trained a classifier on the GPT-labeled dataset. The overall performance of this classifier on the test set is shown in Table 3. It achieved an accuracy of 63%, with a precision of 0.65, recall of 0.62, and F1 score of 0.64. The detailed confusion matrix is presented in Figure 3, which shows the classification behavior across all emotion categories
The confusion matrix shows that all emotion categories are sometimes predicted as Neutral. This is expected, as many Marathi sentences have emotions that are expressed in a very subtle way, making them harder for the model to detect. In such cases, the model often chooses the Neutral label. This problem is common in low-resource languages, where emotions depend more on cultural and contextual clues than on clear emotional words.
There are also some clear patterns of confusion between emotions that are similar in meaning. For example, Pride and Respect are often mixed up. In Marathi, pride is often expressed with respectful language, and respectful statements can sound like pride. Similarly, Happiness and Excitement are confused with each other because they are both positive emotions, with the main difference being the level of intensity. Fear and Surprise are also mixed up, as both can be caused by unexpected events.
In some cases, Sarcasm is classified as Anger/Disgust, which makes sense because sarcasm can carry a tone of irritation or contempt. Sadness is sometimes labeled as Neutral when expressed in a mild way, and as Anger/Disgust when it includes frustration. These patterns show two main challenges: the tendency of the Neutral class to attract unclear cases, and the difficulty of separating emotions that are similar in meaning or context. Better use of context and targeted data augmentation could help improve the model’s performance in these cases.
@@ 4.3 Chain of Translation Prompting (CoTR) vs Non-CoTR Approach
5 Limitations
One limitation of our work is that we used large language models (LLMs) that are mostly trained on English or multilingual data, not specifically on Marathi. Because of this, the models may not fully understand the deeper meanings or cultural context in Marathi sentences.
Another limitation is that our dataset has fewer examples of rare or complex emotions, which makes it harder for the model to learn and predict such emotions correctly. Emotions like Emotional overwhelm and Gentle sorrow are especially difficult to label consistently.
@@ 6 Future Work and Conclusion
In the future, we plan to train LLMs using more Marathi-specific emotion data. This will help the models better understand the language and its emotional tone. We also want to include more sentences that show complex and subtle emotions of Marathi.
We also aim to test our Chain-of-Translation (CoTR) method on more LLMs such as Gemma, Grok, DeepSeek, and Mistral, to see how well it works with other models.